OpenStudy (anonymous):

A water canal has a rectangular cross section 4 meters wide and 3 meters deep. How much water is contained in a 20 meter length of the canal? How much water will be held after 40% has evaporated?

OpenStudy (anonymous):

For part a) This is asking forth volume: V = l*e*h V = (4)(3)(20) V = 240 m^3 for b) if 40% evaporates, you are left with 60% of the water: 60/100* 240 = 144 m^3